### v1

# Towards Two-Stage Counterfactual Learning to Rank

Abstract.
[…]
CLTR aims to learn a ranking policy from user interactions while correcting for the inherent biases in interaction data, such as position bias. Existing counterfactual learning to rank (CLTR) methods assume a single ranking policy that selects top-$K$ ranking from the entire document candidate set. In real-world applications, the candidate document set is on the order of millions, making a single-stage ranking policy impractical. In order to scale to millions of documents, real-world ranking systems are designed in a two-stage fashion, with a candidate generator followed by a ranker. The existing CLTR method for a two-stage offline ranking system only considers the top-1 ranking set-up and only focuses on training the candidate generator, with the ranker fixed. A CLTR method for training both the ranker and candidate generator jointly is missing from the existing literature.
In this paper, we propose a two-stage CLTR estimator that considers the interaction between the two stages and estimates the joint value of the two policies offline. In addition, we propose a novel joint optimization method to train the candidate and ranker policies, respectively. To the best of our knowledge, we are the first to propose a CLTR estimator and learning method for two-stage ranking. Experimental results on a semi-synthetic benchmark demonstrate the effectiveness of the proposed joint CLTR method over baselines.
Learning to rank; Counterfactual learning to rank
submissionid: 16copyright: rightsretainedjournalyear: 2025copyright: ccconference: Proceedings of the 2025 International ACM SIGIR Conference on Innovative Concepts and Theories in Information Retrieval (ICTIR); July 18, 2025; Padua, Italybooktitle: Proceedings of the 2025 International ACM SIGIR Conference onInnovative Concepts and Theories in Information Retrieval (ICTIR) (ICTIR ’25), July 18, 2025, Padua, Italydoi: 10.1145/3731120.3744583isbn: 979-8-4007-1861-8/2025/07ccs: Information systems Learning to rank
1. Introduction
A learning to rank (LTR) method aims to optimize a ranking policy to maximize a given information retrieval (IR) metric (Liu, 2009). Traditionally, LTR policies were trained using manually curated queries and document relevance judgments. It is well known now that manually curating relevance judgments is time consuming, not scalable, and does not always translate to user preferences (Sanderson, 2010; Chapelle and Chang, 2011). As an alternative, LTR from user interactions is scalable because user interactions are cheaper to collect on scale, and user interactions are generally aligned with user interests (Karatzoglou et al., 2013). However, user interactions are noisy and biased indicators of true user preferences, subject to biases such as position bias, and trust bias in search engines (Agarwal et al., 2019b; Joachims et al., 2017).
[…]
Traditionally, CLTR methods assume a single re-ranking policy, i.e., for a given query, the re-ranking policy selects top-$K$ items from the candidate document set, generated via a (lightweight) candidate generator. In a typical CLTR pipeline, the candidate generator is fixed, and only the re-ranker is updated with the CLTR method (Gupta et al., 2024a). To the best of our knowledge, there is no CLTR method that jointly updates both candidate generator and the re-ranker.
In the context of offline contextual bandit learning, Ma et al. (2020) introduced the first two-stage off-policy correction method. The proposed estimator considered the interaction between the candidate generator and the contextual bandit policy. Although effective, the method has the following limitations: (i) it is only designed for the contextual bandit setup (top-1 ranking) and extending it to the CLTR setup is non-trivial since existing CLTR estimators do no consider the interaction between different policies in a cascading setup (Agarwal et al., 2019a; Joachims et al., 2017; Oosterhuis and de Rijke, 2020); and (ii) it only trains the candidate generator using the two-stage estimator while keeping the second-stage bandit policy fixed, typically pre-trained.
We argue that this strategy is not optimal. Initially trained using logged data generated from the production candidate generator, the re-ranker faces potential performance issues when a new candidate generator is deployed with the new candidate policy. This causes a distribution shift in the input data distribution for the re-ranker, potentially leading to an overall degradation in the re-ranker performance when deployed online. To mitigate this, we propose a joint optimization method to optimize the re-ranker and candidate generator via an alternating optimization fashion. Experimental results on a semi-synthetic benchmark demonstrate that the proposed joint optimization method outperforms the two-staged baseline method, where the re-ranker and candidate generator are trained independently. To the best of our knowledge, this work is the first to develop a CLTR estimator and learning method for the two-stage LTR system. We hope our contributions will enable LTR practitioners to apply CLTR methods at a real-world scale.
2. Related Work
[…]
In the context of LTR, Joachims et al. (2017) introduced the first counterfactual learning method to correct for position bias in search engines (Joachims, 2002). They applied the IPS weighting technique, common in the offline contextual bandit literature (Saito and Joachims, 2021). For each user interaction with a document displayed by the ranking system, a click is weighted with the inverse of the document examination probability, a.k.a. document propensity. With the IPS weighting scheme, the effect of the position bias is removed in expectation. As an extension, Oosterhuis and de Rijke (2020) introduced a policy-aware CLTR method for the top-$K$ ranking setup. In a top-$K$ ranking setup, any document ranked beyond the position $K$ gets zero exposure. In this work, we deal with the top-$K$ setting. For a more recent overview of CLTR, we refer to (Gupta et al., 2023a).
2.2. Offline policy learning for contextual bandits
In the context of offline contextual bandits, IPS is commonly used to correct for selection bias over actions introduced by the previously deployed behavior policy (Swaminathan and Joachims, 2015; Swaminathan et al., 2017; Xie et al., 2018; Schnabel et al., 2016; Joachims et al., 2018; Wu and Wang, 2018; Gupta et al., 2023b, 2024b). Briefly, for each (action, context) pair, the corresponding reward is weighted by the ratio of corresponding action probabilities under the new policy and the behavior policy. In expectation, this weighting scheme removes the effect of selection bias introduced by the behavior policy.
2.3. Two-stage offline policy learning
Most work on offline policy learning for contextual bandits and LTR consider a single policy, typically the re-ranking policy, that re-ranks all the candidate documents for a given query context, generated via a candidate generation policy (Saito and Joachims, 2021; Gupta et al., 2023a). Typically, the candidate generation policy is fixed during the off-policy update of the re-ranker, and the interaction between the two policies is generally ignored.
In the context of LTR from manual relevance judgments, Dang et al. (2013) introduced the first joint two-stage ranking method, which considers the interaction between the re-ranker and candidate policy and jointly updates both. Wang et al. (2011) introduced an efficient boosting-based method for joint two-staged ranking.
In offline contextual bandits, Ma et al. (2020) proposed a two-stage off-policy learning objective that considers the interaction between the first and second-stage contextual bandit policies. They propose a policy learning method to train the candidate generator considering the response of the re-ranker on the candidate generated at the first stage.
3. Background
[…]
For a given query $q$ ($q\sim P(Q)$), a ranked list $y=\{d_{1},d_{2},\ldots,d_{k}\}$ generated by a ranker $\pi$, we define the utility of the list $y$ as:
(1) $U(y\mid q,\pi)=\sum_{d\in D}\alpha_{k(d)}P(R=1\mid q,d)=\sum_{d\in D}\alpha_{k(d)}R_{q,d},$
where $k(d)$ is the rank of the document $d$ in the list $y$, and $\alpha_{k(d)}$ is the weight for the document $d$ at the rank $k(d)$, and $P(R=1\mid q,d)=R_{q,d}$ is the probability of relevance for the query, document pair $(q,d)$. Given the ranking utility, the goal of a LTR algorithm is to train a ranking policy $\pi$ to optimize the following overall utility function:
(2) $\begin{split}U(\pi)&=\mathbb{E}_{q\sim P(Q)}\mathbb{E}_{y\sim\pi(\cdot\mid q)}\left[U(y\mid q,\pi)\right]\\ &=\mathbb{E}_{q}\left[\sum_{d\in D}\rho(d\mid q,\pi)R_{q,d}\right],\end{split}$
where the choice of the document weight $\rho(d\mid q,\pi)=\mathbb{E}_{y\sim\pi}\!\left[\alpha_{k(d)}\right]$ (Eq. 1) defines the IR metric that is being optimized as a result. For example, the choice of $\alpha_{k(d)}=(\log_{2}(\textrm{rank}(d\mid y)+1))^{-1}$ means that we are optimizing for the NDCG metric (Liu, 2009).
Traditional LTR methods assume access to the true manually graded document relevance scores: $R_{q,d}$ (Eq. 1). However, manually curating relevance judgments in practice is not scalable, and relevance judgments are generally not aligned with user interests (Chapelle and Chang, 2011; Qin et al., 2010). As an alternative, CLTR considers the user interactions directly. CLTR assumes access to the click log dataset $\mathcal{D}$ from a logging policy ($\pi_{0}$), $\mathcal{D}=\big{\{}q_{i},y_{i},c_{i}\big{\}}^{N}_{i=1},$ with user issued queries $q_{i}$, the ranked list from the ranking policy $y_{i}$, and, finally, user clicks ($c_{i}\in{0,1}$) on the list $y_{i}$ indicate whether the document was clicked or not. In this work, we assume that user clicks follow the examination hypothesis (Chuklin et al., 2015; Joachims et al., 2017):
(3) $P(C=1\mid d,q,k)=P(E=1\mid k)P(R=1\mid d,q),$
which states that for a given query $q$, the probability of clicking on a document $d$ depends on the probability of examination at a given rank $k$ of the document and its relevance.
Given the click model (Eq. 3), to define the IPS-based CLTR objective, following Oosterhuis and de Rijke (2020), we define the document propensity under a policy $\pi$ as:
(4) $\rho(d\mid q,\pi)=\mathbb{E}_{y\sim\pi}\left[P(E=1\mid k(d))\right].$
A similar propensity definition was used in other CLTR work (Yadav et al., 2021; Gupta et al., 2023c; Oosterhuis and de Rijke, 2021a). The choice of rank-based examination probability as the propensity sets the overall utility function as the expected number of clicks (Eq. 2). Finally, the CLTR utility function is defined as:
(5) $\hat{U}(\pi)=\frac{1}{N}\sum_{i=1}^{N}\sum_{d\in D}\frac{\rho(d)}{\rho_{0}(d)}c_{i}(d),$
with $\rho_{0}(d)$ as the propensity score (Eq. 4) for the logging policy $\pi_{0}$.
3.2. Two-stage off-policy learning
In the context of offline contextual bandits, Ma et al. (2020) proposed a two-stage off-policy learning method, where the objective function is defined with respect to the bandit feedback dataset:
(6) $\mathcal{D}=\big{\{}s_{i},a_{i},r_{i}\big{\}}^{N}_{i=1},$
where $s_{i}$ is the context/state, $a_{i}\sim\pi(\cdot\mid s_{i})$ is the action sampled from the policy $\pi$, and $r_{i}$ is the reward for the (state, action) pair. The policy $\pi$ is defined as the mixture of the candidate policy which scores a size $k$ list: $A^{k}$, $p(A^{k}\mid s)$ and the re-ranker $q(a\mid A^{k},s)$: $\pi(a\mid s)=\sum_{A^{k}}p(A^{k}\mid s)q(a\mid A^{k},s)$. Given this, the two-stage off-policy objective is defined as:
(7) $\hat{U}(\pi)=\frac{1}{N}\sum_{i=1}^{N}\frac{\pi(a_{i}\mid s_{i})}{\pi_{0}(a_{i}\mid s_{i})}r_{i}.$
In their proposed method, only the candidate policy $p(A^{k}\mid s)$ is updated via a stochastic gradient method; the re-ranker $q(a\mid A^{k},s)$ is kept fixed.
4. Method: Two-stage Counterfactual Learning to Rank
This section introduces our main contribution: a novel two-stage CLTR objective and a joint learning method.
4.1. Two-stage LTR
We first define the two-stage LTR objective with relevance labels. Given a candidate generator policy $\pi_{c}$, the candidate list $y_{c}\sim\pi_{c}$ of length $K_{2}$, the re-ranker policy $\pi_{r}$, and the final ranked list displayed to the user $y_{r}\sim\pi_{r}$ of length $K$, similar to the utility function for a given list in a single-stage setup $U(y\mid q,\pi)$ (Eq. 1), the two-stage LTR objective is given by:
$U(\pi_{c},\pi_{r})$ $=\mathbb{E}_{q\sim P(Q)}\mathbb{E}_{y_{c}\sim\pi_{c}(\cdot\mid q)}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[U(y_{r}\mid q,\pi)\right]$
$=\mathbb{E}_{q\sim P(Q)}\mathbb{E}_{y_{c}\sim\pi_{c}(\cdot\mid q)}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[\sum_{d\in D}\alpha_{k(d)}R_{q,d}\right]$
(8) $=\mathbb{E}_{q\sim P(Q)}\left[\sum_{d\in D}\rho_{c,r}(d)R_{q,d}\right],$
where $\rho_{c,r}(d)=\mathbb{E}_{y_{c}\sim\pi_{c}(\cdot\mid q)}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[\alpha_{k(d)}\right]$ is the new document weight, equivalent to the single-stage case (Eq. 2).
4.2. Method: A novel two-stage CLTR
Given the true joint utility function for the two-stage ranking with the true relevance (Eq. 8), we now define the CLTR objective with the logged data $\mathcal{D}$ (Eq. 6):
(9) $\hat{U}(\pi_{r},\pi_{c})=\frac{1}{N}\sum_{i=1}^{N}\sum_{d\in D}\frac{\rho_{r,c}(d)}{\rho_{0}(d)}c_{i}(d).$
Similarly to the single-stage CLTR objective (Eq. 5), we set the weight $\alpha_{k(d)}=P(E=1\mid k(d))$. As a result, the overall objective optimizes for expected clicks on relevant documents (Oosterhuis, 2023). To the best of our knowledge, we are the first to introduce a two-staged CLTR objective for real-world ranking systems.
Theorem 4.1.
The counterfactual objective for the two stage CLTR method (Eq. 9) is unbiased in expectation, i.e.,
(10) $\mathbb{E}_{q,y\sim\pi_{0},c}\left[\hat{U}(\pi_{r},\pi_{c})\right]=U(\pi_{r},\pi_{c}).$
Proof.
The expected value of the two-stage objective is given by:
$\mathbb{E}_{q,y\sim\pi_{0},c}\left[\hat{U}(\pi_{r},\pi_{c})\right]$ $=\mathbb{E}_{q}\left[\sum_{d\in D}\frac{\rho_{r,c}(d)}{\rho_{0}(d)}\mathbb{E}_{y\sim\pi_{0},c}\left[c(d)\right]\right]$
(11) $=\mathbb{E}_{q}\left[\sum_{d\in D}\rho_{r,c}(d)R_{q,d}\right],$
where we apply the click model (Eq. 3) in the first step. ∎
Intuitively, to maximize the objective in case of a click on a given document $d$, the optimizer will push the expected document weight under both policies $\rho_{r,c}(d)$ to the maximum. To maximize the expected weight, the candidate generator will push the document higher in the candidate list, which gives the re-ranker a chance to finally display the document to the user at a higher rank.
The candidate generator can be selected as a lightweight model for efficiency, generating the candidate list quickly. The re-ranker can be a complex model that incorporates additional features to improve the precision of the results. However, for simplicity, we use the same model for both candidate generator and re-ranker and leave the effects of different architecture choices for future work.
4.3. Joint optimization for two-stage CLTR
For optimization, we use stochastic gradient descent. To estimate the gradient with respect to the policy $\pi$, in this work, we make use of the general log-derivate trick of the REINFORCE algorithm (Williams, 1992), with the gradient of the metric weight term following:
(12) $\nabla_{\!\pi}\rho(d)=\mathbb{E}_{y\sim\pi}\big{[}P(E=1\mid k)\nabla_{\!\pi}\log\pi(y\mid q)\big{]}.$
Ma et al. (2020) proposed to optimize the candidate generator policy while keeping the re-ranking policy constant during the optimization (the re-ranker is pre-trained on the same logged data). For the CLTR context, this translates into estimating the gradient:
(13) $\nabla_{\!\pi_{c}}\hat{U}(\pi_{r},\pi_{c})=\frac{1}{N}\sum_{i=1}^{N}\sum_{d\in D}\frac{\nabla_{\!\pi_{c}}\rho_{r,c}(d)}{\rho_{0}(d)}c_{i}(d).$
While this objective considers the feedback from the re-ranker for training the candidate generator, this strategy is overall sub-optimal. The re-ranker policy is pre-trained on the log data generated by the logging policy $\pi_{0}$. This means that it is optimized with respect to candidates generated by the logging policy during training. At inference time, if the candidate generator is switched to a different one, its performance might suffer because of the distribution shift.
For two-stage optimization, if the re-ranker is kept fixed and the candidate is updated during training, then at test time, the distribution of the candidates will change, and the re-ranker might underperform. To remedy this problem, we propose a joint optimization method, where both the candidate and re-ranker are updated during training. We follow an alternating optimization method, wherein for each minibatch, the candidate generator and the re-ranker are updated alternatingly. This ensures that the re-ranker considers the candidate generator’s feedback during training.
Table 1. NDCG@10 performance of different two-stage CLTR methods with varying candidate list sizes ($K_{2}$) and varying amounts of click logs ($N$). The numbers reported are averages over 25 independent runs.
$K_{2}=500$ $K_{2}=1,000$ $K_{2}=1,500$
Method $N=0.1M$ $N=0.32M$ $N=1M$ $N=0.1M$ $N=0.32M$ $N=1M$ $N=0.1M$ $N=0.32M$ $N=1M$
Baseline 0.404 (0.000) 0.437 (0.001) 0.470 (0.001) 0.401 (0.001) 0.440 (0.001) 0.469 (0.001) 0.400 (0.001) 0.441 (0.001) 0.474 (0.002)
Indepen. 0.410 (0.001) 0.454 (0.001) 0.489 (0.001) 0.416 (0.001) 0.465 (0.015) 0.496 (0.001) 0.416 (0.001) 0.455 (0.001) 0.496 (0.001)
Joint opt. 0.420 (0.001) 0.470 (0.001) 0.504 (0.001) 0.426 (0.001) 0.479 (0.001) 0.504 (0.001) 0.424 (0.001) 0.472 (0.001) 0.510 (0.001)
Gradient calculation. As the choice of the ranking policies in the candidate generator and re-ranker, we choose a Plackett-Luce model, similar to previous work in CLTR (Oosterhuis, 2021, 2023; Gupta et al., 2023c). For optimization, we use stochastic gradient descent. For the gradient of the document weight with respect to the candidate policy, we apply the log-derivative/REINFORCE trick (Williams, 1992; Yadav et al., 2021; Gupta et al., 2023c). The gradient expression is given as:
(14) $\begin{split}&\nabla_{\!\pi_{c}}\rho_{r,c}(d)={}\\ &\mbox{}\hskip 8.53581pt\mathbb{E}_{y_{c}\sim\pi_{c}}\big{[}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[P(E=1\mid k)\right]\nabla_{\!\pi}\log\pi_{c}(y_{c}\mid q)\big{]}.\end{split}$
Similarly, the gradient of the document weight with respect to the re-ranker policy can be expressed as:
(15) $\begin{split}\mbox{}&\nabla_{\!\pi_{r}}\rho_{r,c}(d)={}\\ \mbox{}&\mbox{}\hskip 5.69054pt\mathbb{E}_{y_{c}\sim\pi_{c}}\big{[}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[P(E=1\mid k)\nabla_{\!\pi}\log\pi_{r}(y_{r}\mid y_{c},q)\right]\big{]}.\end{split}$
5. Experimental Setup
For our experiments, we follow the semi-synthetic experimental setup, common in the CLTR literature (Joachims et al., 2017; Oosterhuis and de Rijke, 2020; Gupta et al., 2023c, 2024c). The standard LTR datasets involve a query and a pre-selected document list with the corresponding relevance judgments, for example, the MSLR30K dataset (Qin et al., 2010). The logging policy is trained on 3% of the queries to simulate a production ranker, which is used to generate the logged dataset following a position-based user model (Craswell et al., 2008; Oosterhuis and de Rijke, 2020; Joachims et al., 2017). In the LTR datasets, documents are already prefiltered from the entire document pool and in the order of $\mathbf{O}(100)$, which is not reflective of a real-world setup, where the candidate document set is of much larger order. As a result, the LTR datasets are unsuitable for a realistic large-scale simulation of a two-staged CLTR system. Instead, we rely on the MovieLens-1M recommender systems dataset for the simulation in this work. The dataset consists of user ratings of items. Following previous work (Zou et al., 2020; Zhang et al., 2019; Zhao et al., 2013), we assume that the ratings reflect the true user intent and are not biased by the recommendations to the user. We assume a rating value of $>$ 3 as a positive reward and a zero otherwise. We randomly sample $10\%$ of users with the true rating as the ground truth for policy evaluation. This simulates a real-world setting where a production policy is trained on a small fraction of the ideal relevance labels, then deployed to collect click signals (Joachims et al., 2017; Agarwal et al., 2019a). We simulate a top-$K$ ranking setup (Oosterhuis and de Rijke, 2020; Gupta et al., 2023c) where any item beyond the top-$K$ slot gets no exposure. The clicks are simulated via a position-based click model (Craswell et al., 2008; Joachims et al., 2017), where we define the examination probability as follows:
(16) $P(E=1\mid q,d,y)=\begin{cases}\textrm{rank}(d|y)^{-1}&\text{if }\textrm{rank}(d\mid y)\leq 10,\\ 0&\text{otherwise}.\end{cases}$
For our experiments, we define the query $q$ as the user context, and an item is a document $d$. Following previous work (Oosterhuis and de Rijke, 2021b, a; Oosterhuis, 2023; Gupta et al., 2023c), we use frequency estimates for the propensity $\rho_{0}(d)$ (Eq. 13). We use a simple matrix factorization model for the re-ranker and candidate generator policy, initialized with user and item embeddings of dimension 50, generated via applying SVD on the initial user-item rating matrix. Exploring the effect of different model architectures is out of scope and will be in future work. For estimating the policy gradient expectations (Eq. 14, 15), we use 300 Monte-Carlo samples (Yadav et al., 2021; Oosterhuis, 2023; Gupta et al., 2023c). For optimization, we use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $0.01$ for all methods. Finally, the following methods are included in our comparisons:
  1. (1)
    Baseline. As a baseline, we follow the setup from previous work (Ma et al., 2020), wherein we pre-train a re-ranker on the click data and only train the candidate generator using the two-stage objective (Eq. 9) and evaluate via the two-stage LTR objective (Eq. 8).
  2. (2)
    Independent. We train the candidate generator and reranker independently and evaluate via the two-stage objective (Eq. 8).
  3. (3)
    Joint optimization. Our proposed method (Sect. 4.3) is where we jointly optimize the candidate generator and re-ranker in an alternating optimization fashion.
6. Results
Table 1 reports the NDCG@10 results for different methods. We report the results with varying the candidate list sizes $K_{2}\in\{500$, $1000$, 1500}1500\}1500 } for realistic real-world evaluation. The table shows that the joint optimization method (Sect. 4.3) performs the best compared to the other methods.
The baseline method from the two-stage bandit work (Ma et al., 2020), where the re-ranker is pretrained from the data generated from the logging candidate policy, performs worst among all methods, confirming our initial hypothesis that changing the candidate ranking distribution from training to inference can result in an overall performance drop for the re-ranker policy. The proposed joint optimization method for the candidate generator and the re-ranker performs the best across different candidate list sizes and logged data. This shows that the joint optimization strategy is optimal for the two-stage CLTR methods. The performance of each method improves with increasing logged data size ($N=0.1M,0.32M,1M$), consistent with previous work (Oosterhuis and de Rijke, 2020; Joachims et al., 2017; Gupta et al., 2023c; Jagerman et al., 2020; Oosterhuis and de Rijke, 2021b).
Note that in this work, we use the same architecture for the candidate generator and the re-ranker, i.e., a matrix factorization model. Exploring the effects of different neural architectures for the candidate generator and re-ranker is beyond the scope of this work, as our focus is on introducing joint optimization for the two-stage CLTR. We will study it in future work.
7. Conclusion
In this work, we have studied two-stage ranking systems from a CLTR perspective. This is an important problem that has not been adequately discussed in the contemporary literature of CLTR; in this work, we address this gap. First, we propose a novel single LTR two-stage objective function (Eq. 8) for optimizing LTR systems with access to the ideal relevance judgments. Next, we propose a two-stage CLTR objective (Eq. 9) that accounts for both the candidate generator and re-ranker jointly. To optimize the overall system, we propose a joint optimization strategy. We also provide the gradients for first-order optimization methods.
To the best of our knowledge, we are the first to tackle the problem of two-stage LTR systems and a counterfactual learning and evaluation method for two-staged systems. As part of future work, we wish to extend the formulation from a contextual bandit setup to a reinforcement learning setup (Gao et al., 2023).
Acknowledgements.
This research was supported by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union’s Horizon Europe program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.

### v2

@@
 The gradient expression is given as:
-(14) $\begin{split}&\nabla_{\!\pi_{c}}\rho_{r,c}(d)={}\\ &\mbox{}\hskip 8.53581pt\mathbb{E}_{y_{c}\sim\pi_{c}}\big{[}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[P(E=1\mid k)\right]\nabla_{\!\pi}\log\pi_{c}(y_{c}\mid q)\big{]}.\end{split}$
+(14) $\begin{split}&\nabla_{\!\pi_{c}}\rho_{r,c}(d)={}\\ &\mbox{}\hskip 8.53581pt\mathbb{E}_{y_{c}\sim\pi_{c}}\big{[}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[P(E=1\mid k)\right]\nabla_{\!\pi_{c}}\log\pi_{c}(y_{c}\mid q)\big{]}.\end{split}$
 Similarly, the gradient of the document weight with respect to the re-ranker policy can be expressed as:
-(15) $\begin{split}\mbox{}&\nabla_{\!\pi_{r}}\rho_{r,c}(d)={}\\ \mbox{}&\mbox{}\hskip 5.69054pt\mathbb{E}_{y_{c}\sim\pi_{c}}\big{[}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[P(E=1\mid k)\nabla_{\!\pi}\log\pi_{r}(y_{r}\mid y_{c},q)\right]\big{]}.\end{split}$
+(15) $\begin{split}\mbox{}&\nabla_{\!\pi_{r}}\rho_{r,c}(d)={}\\ \mbox{}&\mbox{}\hskip 5.69054pt\mathbb{E}_{y_{c}\sim\pi_{c}}\big{[}\mathbb{E}_{y_{r}\sim\pi_{r}(\cdot\mid y_{c},q)}\left[P(E=1\mid k)\nabla_{\!\pi_{r}}\log\pi_{r}(y_{r}\mid y_{c},q)\right]\big{]}.\end{split}$
 5. Experimental Setup